### v2

# MOON: Multi-Objective Optimization-Driven Object-Goal Navigation Using a Variable-Horizon Set-Orienteering Planner

Abstract
Object-goal navigation (ON) enables autonomous robots to locate and reach user-specified objects in previously unknown environments, offering promising applications in domains such as assistive care and disaster response. Existing ON methods—including training-free approaches, reinforcement learning, and zero-shot planners—generally depend on active exploration to identify landmark objects (e.g., kitchens or desks), followed by navigation toward semantically related targets (e.g., a specific mug). However, these methods often lack strategic planning and do not adequately address trade-offs among multiple objectives. To overcome these challenges, we propose a novel framework that formulates ON as a multi-objective optimization problem (MOO), balancing frontier-based knowledge exploration with knowledge exploitation over previously observed landmarks; we call this framework MOON (MOO-driven ON). We implement a prototype MOON system that integrates three key components: (1) building on QOM [IROS05], a classical ON system that compactly and discriminatively encodes landmarks based on their semantic relevance to the target; (2) integrating StructNav [RSS23], a recently proposed training-free planner, to enhance the navigation pipeline; and (3) introducing a variable-horizon set orienteering problem formulation to enable global optimization over both exploration and exploitation strategies. This work represents an important first step toward developing globally optimized, next-generation object-goal navigation systems.
I Introduction
Object-goal navigation (ON) is a fundamental task for autonomous mobile robots, where the robot must efficiently locate and reach a target object specified by user instructions, typically provided in the form of an image [1] or natural language [2], within an unknown environment. ON technology holds promise for various real-world applications, including autonomous assistive robots in caregiving and welfare settings to improve the quality of life for the elderly and people with disabilities [3], automation and labor-saving in logistics and warehouse management through efficient picking and inventory handling [4], and operations in safety-critical situations such as disaster response and search and rescue missions [5]. Achieving ON requires a wide range of technologies, including object detection, image classification, semantic segmentation, feature matching, natural language processing, multimodal understanding, SLAM, planning, reinforcement learning, collision avoidance, imitation learning, user interface design, and simulation environment and benchmark development. As such, ON has recently gained status as a standard problem spanning robotics, AI, computer vision, and computer science.
Figure 1: ON setup. This study explicitly models two types of visual recognition modules: a long-range landmark detector and a short-range target detector. When the sensing range of the landmark detector substantially exceeds that of the target detector, the landmark visiting order optimization problem (LOOP) becomes a key factor in efficient navigation. Specifically, an observation path exemplified by plan B, which maximizes the information gain-to-cost ratio, is preferable to those represented by plan A or plan C.
Various approaches to ON have been explored, including training-free planners (TFP) [1], reinforcement learning-based planners (RLP) [6] that learn strategies through interaction with the environment, and zero-shot planners (ZSP) [2] that generalize to novel goals based on pretraining. For example, StructNav [1], a novel TFP implementation, prioritizes exploratory behaviors in the early stage and switches to goal-directed landmark-based navigation in the later stage, where ”landmark objects” are those semantically related to the target. Unlike many learning-based methods that suffer from high training costs and lack of generalization or interpretability, StructNav adopts a modular architecture. It combines structured scene representations using 2D occupancy maps, semantic point clouds, and spatial scene graphs with semantic frontier exploration guided by large language models and scene statistics. This results in explainable, robust navigation without training costs and demonstrates superior performance over existing learning-based methods, with strong potential for real-world deployment.
Despite these advances, we observe that efforts to leverage the long-standing knowledge of the optimization techniques community—especially optimal planning techniques [7]—in ON methods are either overlooked or remain underdeveloped (Fig. 1). Consequently, many systems fall into myopic behaviors and fail to balance exploratory and goal-directed actions effectively. While existing ON methods often encourage visiting semantically relevant landmarks, they largely neglect strategic planning of the visiting order and global path optimization, thereby underutilizing valuable insights from the optimization community. This issue becomes critical in real-world environments, which are often large-scale and complex, yet many existing methods rely on idealized simulators such as Habitat-Sim [8] and AI2-THOR [9], which oversimplify sensor models and spatial complexity, limiting real-world applicability. Furthermore, in multi-robot scenarios, landmark maps can be shared to build a richer global knowledge base, but effective utilization of such shared information for coordinated strategic navigation remains underexplored. We argue that integrating intuitive, myopic high-level reasoning from existing TFP/RLP/ZSP frameworks with globally optimal planning is essential for effective and efficient ON in large, complex real-world settings.
In this work, we reformulate the standard single-target ON task as a multi-objective optimization-driven object goal navigation (MOON), and as a case study, model two conflicting objectives: frontier exploration for discovering unknown landmarks and efficient touring of previously observed landmarks (Fig. 2)111 The ”moon and star” representation was introduced as a schematic or cartoon-like abstraction to emphasize visual clarity while maintaining mathematical tractability. Nevertheless, our framework is not limited to point landmarks or circular sensing ranges and can be generalized to a wide range of application domains. . Our formulation emphasizes the intrinsic multi-objective nature of ON tasks and is both orthogonal and complementary to existing problem settings, such as multi-ON tasks [10], which address multiple target objects in an optionally specified order, and general multi-objective navigation [11]. Specifically, we model ON planning as a variable-horizon optimization problem aiming to maximize average SPL (Success weighted by Path Length [12]) based on incrementally constructed landmark maps. Although our method is compatible with the TFP, RLP, and ZSP frameworks, we adopt the TFP framework from [1] as the basis for our experiments. In this framework, the robot detects objects from RGBD and semantic images and semantically estimates landmark objects related to the target. We then extend TFP by introducing a query-based occupancy map (QOM) [13], which compactly describes essential landmark information by discarding data irrelevant to object-target navigation and focusing solely on relevant landmarks. Finally, we formulate the landmark visiting order optimization problem (LOOP) as a set orienteering problem (SOP) and further extend it to a variable-horizon SOP that incorporates map updates. Notably, the proposed method facilitates map sharing and multi-objective navigation, enhancing the efficiency and scalability of ON tasks by integrating semantic reasoning with optimal planning.
Figure 2: Two conflicting objectives. The crescent moon-shaped yellow region represents the area newly observed by the robot. Landmarks are depicted as stars, with the previously observed area detected by the landmark detector shown in white, and the unobserved area shown in black. While unvisited landmarks exist within the previously mapped region (white), potential unvisited landmarks may also be present in the unmapped region (black). Determining which region to prioritize constitutes an ill-posed and computationally challenging planning problem.
I-A Relation to Prior Work
The current study is related to several prior works on object-goal navigation (ON) and multi-objective optimization (MOO) problems.
An exploratory strategy that utilizes a wide-view image as a landmark map for exploration task was first proposed in [14]. However, this work focused solely on an alternative exploration task—namely, floor cleaning—and did not address the object-goal navigation problem itself.
The QOM-based object-goal navigation (ON) approach employed in the current research is not new and was developed around 2005 by researchers at the Mathematical Engineering Laboratory, Department of Mechanical Engineering, Graduate School of Kyushu University [13] (QOM [IROS05]). At that time, an active vision task that was conceptually equivalent to ON was studied. However, unlike the now-common ON setup, which involves exploration of the physical real-world environment, the robot’s workspace in that earlier research was its internal memory space—namely, visual experiences—rather than the physical environment itself, a context in which offline reinforcement learning [15] plays a critical role. Several follow-up studies have since been conducted.
In the field of MOO, the same laboratory has also made longstanding contributions across a wide range of topics, including transport systems, ant colony optimization, passenger railway vehicle operation planning, genetic algorithms, Type-1 servo systems, vibration control system design, behavior acquisition of mobile robots, edging control, and tail crash control [16].
In recent years, we have independently developed a series of methods aligned with the modern ON setup, including TFP [17] (CON [CoRR24a]), RLP [18] (ALCON [CoRR24b]), and ZSP [19] (LGR [CoRR25]).
However, to the best of our knowledge, no prior work has integrated object-goal navigation (ON) with multi-objective optimization (MOO) in a way that is applicable to wide-area ON missions.
a
b
Figure 3: Prototype system. Our prototype is based on StructNav [RSS23] [1], with modifications limited to the essential components of our proposed method. Figure 1(a) illustrates StructNav, whereas Figure 1(b) depicts our approach. The key differences between the two systems are highlighted in red.
II System Overview
This section describes our method that oversees the myopic TFP/RLP/ZSP in conventional ON systems by employing a multi-objective optimization planner that guarantees global optimality (Fig. 3). We begin by introducing our extended problem setting, which leverages a variable-horizon incremental landmark map (VHIM), to clarify the motivation behind our approach (II-A). We then present the system design to address this problem. Specifically, we first review the TFP planner proposed in [1] (II-B), then explain how it is extended to a QOM-based planner following the ALCON framework [18] (II-C), and finally, we introduce our proposed multi-objective optimization planner (II-D).
Figure 4: An oblique view from a high vantage point serves as prior information for the exploration task. The oblique view depicted at the top was employed in an alternative exploration task, namely floor cleaning [IROS01], originally presented in [14]. Over 20 years ago, when the original study was conducted, this view image exceeded the capabilities of the image recognition systems available at that time. Consequently, the approach employed a procedure whereby the robot autonomously selected the widest-area view from its visual experience and transmitted it to a teleoperator. The teleoperator subsequently used this view to plan the exploration, with an example of the planned path indicated by the white line in the figure. The middle part illustrates how the image, combined with a textual prompt, was used as input to a recent large language model (LLM). The bottom part shows the response from the LLM. The results indicate that the LLM demonstrates a level of scene understanding sufficient to provide prior knowledge for robotic path planning. The primary remaining challenge pertains to exploration path planning, which is the focus of the present study.
Figure 5: Variable-horizon incremental map (VHIM). This figure presents three additional examples. In each image, the robot’s workspace of interest is indicated by a colored ellipse. The top image shows a prior map of a previously visited location. The middle image depicts an aerial view acquired by a teammate drone. The bottom image illustrates a wide-angle view captured prior to entering the workspace, specifically at an intersection. These images correspond to those presented in [20], [21], and [22], respectively.
II-A Motivation: Variable-Horizon Incremental Map (VHIM)
To simplify the discussion, we consider a typical ON strategy in which the agent efficiently explores a workspace by detecting landmark objects (e.g., kitchen, desk) that are semantically related to the target object (e.g., a specific mug) and uses these landmarks as cues for locating the target. This example serves to illustrate the motivation behind our work.
With recent advances in ON and AI technologies, there is a growing demand to extend ON systems using variable-horizon incremental landmark maps (VHIMs) that allow the mapping horizon to dynamically change during incremental mapping (Fig. 4), in contrast to the fixed or small-scale maps typically used in conventional ON tasks. In particular, the following scenarios highlight the emerging significance of such maps:
  1. 1.
    High-elevation view images that overlook the entire workspace are becoming practical as landmark maps for ON [14] (Fig. 4).
  2. 2.
    In multi-objective navigation scenarios, pre-built maps can serve directly as a prior landmark map during ON tasks [11] (Fig. 5a).
  3. 3.
    In multi-robot coordination scenarios, aerial images captured by a partner drone serve as an alternative landmark map for ON [21] (Fig. 5b).
  4. 4.
    A wide-area image taken from a distant viewpoint before the robot enters the workspace is also utilized as a landmark map [22] (Fig. 5c).
While expanding the workspace significantly increases uncertainty in the frontier exploration problem (FEP) in unknown areas, it also substantially raises the computational complexity of the landmark visiting order optimization problem (LOOP) in known areas. Recognizing this challenge serves as the foundation for our multi-objective optimization approach, which aims to achieve global optimality.
II-B Planner Architecture in StructNav
As previously mentioned, our prototype system is built upon the TFP planner from the StructNav framework [1], which comprises the following three stages:
Stage 1: Structured scene representation. Semantic segmentation (Mask R-CNN) is used to generate a semantic category image $I^{sem}_{t}$ from the RGB input. Simultaneously, Visual SLAM (RTAB-Map) estimates the current position $\hat{s}_{t}^{l}$ and updates the RGB reconstruction $P^{rgb}_{t}$. A 2D occupancy map $M_{t}$ is generated based on the reconstructed image, and a 3D semantic point cloud $P^{sem}_{t}$ is created using depth and semantic information. Finally, a spatial scene graph $G_{t}=(V_{t},E_{t})$ is built from the point cloud.
Stage 2: Scene graph construction. The scene representation consists of three elements: the occupancy map $M_{t}$, semantic point cloud $P^{sem}_{t}\in\mathbb{R}^{k_{t}\times 4}$, and spatial scene graph $G_{t}=(V_{t},E_{t})$. The occupancy map is dynamically extended via SLAM; the point cloud includes each point’s $(x,y,z)$ coordinates and semantic labels. The graph’s nodes $V_{t}$ represent object-centric keypoints, and edges $E_{t}$ represent relative positions between objects.
Stage 3: Goal and action selection. A set of frontier cluster centers $F_{t}=\{f^{t}_{0},...,f^{t}_{N}\}$ is extracted at the boundary between explored and unexplored regions. The Utility module selects the most promising intermediate goal $x^{goal}_{t}$ using the frontier candidates and the scene graph $G_{t}$. A Global/Local Planner then determines the action $a_{t}$ to control the agent.
Figure 6: Query-based Occupancy Map (QOM). The top image shows a QOM constructed by an ON system and is reproduced from [13]. The middle image illustrates the exploration results of the ON task, visualized as separating boundaries in both the feature space and the real-world configuration space. The bottom image depicts real-world target objects used as inputs to the QOM construction process. The middle and bottom images are adapted from [23].
II-C QOM-based ALCON Planner
The Query-Based Occupancy Map (QOM) [13] extends the conventional concept of occupancy maps [24] from representing the existence probability of objects in each grid cell to representing the existence probability of a specific target object (Fig. 6). In the original formulation, QOM leveraged the advanced visual recognition capabilities of a remote human operator to estimate the likelihood of target presence, by training and repeatedly employing a machine learning model (e.g., SVM) that imitates human operators. In contrast, the present study replaces human vision with modern computer vision techniques, enabling autonomous inference.
The QOM addresses the trade-off between expressiveness and compactness that has long challenged existing methods. In traditional high-dimensional feature maps, each cell or point in an occupancy grid or point cloud is represented using high-dimensional embedding vectors based on appearance, geometry, and semantics. While these representations offer high fidelity, they suffer from poor compactness. Conversely, bag-of-words models provide lightweight descriptions through quantization but lack representational richness.
QOM achieves a favorable trade-off by focusing solely on target objects. Specifically, it encodes each cell with a compact similarity-based representation relative to the target object. This approach allows for efficient memory usage and fast inference, while maintaining sufficient discriminative performance for practical applications.
Recently, QOM was employed in [17] as a compact yet informative grid map that records the probability of the target object’s presence, and in the ALCON framework in [18], QOM was further utilized to construct a scalable navigation route graph. Specifically, the spatial scene graph $G_{t}$ used in StructNav is replaced in ALCON with a globally optimal route graph. Path planning is then performed on this graph using an actor-critic reinforcement learning (RL) planner. In this context, a natural implementation of the QOM involves evaluating how relevant each cluster in the semantic point cloud is to the target object. Following StructNav [1], this relevance can be estimated using BERT embeddings and cosine similarity. The proposed ALCON framework [18], while building on the foundation of StructNav, is notable for replacing the myopic frontier-based planner in StructNav with an RL-based planner.
II-D Extension in the MOON framework
In the current study, we introduce a solver for the variable-horizon set-orienteering problem (VH-SOP) that has the competence to supervise previously developed systems—TFP (Training-Free Planner) [17], RLP (Reinforcement Learning-based Planner) [18], and ZSP (Zero-Shot Planner) [19]—with the goal of achieving globally-optimal path-planning.
Assuming that an object-goal navigation (ON) planner—such as TFP, RLP, or ZSP—has identified regions within the workspace likely to contain landmark objects and aggregated them into a Query-based Occupancy Map (QOM), our system operates as follows. First, it extracts landmarks that are highly relevant to the target object. Next, it computes a set of candidate viewpoints $V$ from which these landmarks can be observed. Finally, it constructs a navigation graph $(V,E)$ based on the QOM, as described in a later section.
To formulate the Set Orienteering Problem (SOP), we explicitly distinguish between two types of visual recognition models: a long-range landmark detector with a sensing range $L$ (e.g., $L=100$ m), and a short-range target detector with a sensing range $R$ (e.g., $R=3$ m). The long-range detector can identify semantically relevant landmarks (e.g., kitchen, desk) within its wide sensing range $L$, and as $L$ increases, the number of candidate landmarks also grows, thereby escalating the complexity of planning. In contrast, the short-range detector is capable of recognizing the target object (e.g., a specific mug) only within its limited range $R$, which serves as a crucial constraint in determining whether a given viewpoint qualifies as a valid landmark visit.
We model the navigation graph such that nodes $V$ represent viewpoints from which landmarks can be observed, and edges $E$ represent traversable paths between node pairs with associated navigation costs. Each node is also annotated with the ID of the observable landmark. A group of nodes that share the same landmark ID is referred to as a node cluster. Visiting multiple nodes within the same cluster yields negligible or no additional informational gain compared to visiting only one of them. This assumption is natural considering the relationship between landmarks and viewpoints. For each node cluster, we estimate the utility of observing the associated landmark and assign the utility score to all nodes within the cluster. Additionally, for every node pair, we assess reachability and establish edges between reachable nodes, assigning edge costs accordingly. All edge costs can be computed by executing Dijkstra’s algorithm once per node.
Finally, we determine a path that maximally utilizes new knowledge under a variable horizon plan using the SOP solver. Details of this final procedure are provided in Section III-A.
II-E Simplification Assumption
In this study, we assume that the robot’s self-location is known. For cases where the location is unknown, please refer to the multi-robot continual ON (CON) approach in [17].
III A MOON Planner
In this section, we formulate the object navigation (ON) as a multi-objective optimization problem, and present one specific solution.
A multi-objective optimization problem is generally formulated as:
$\text{minimize}\quad\mathbf{f}(\mathbf{x})=\left(f_{1}(\mathbf{x}),f_{2}(\mathbf{x}),\dots,f_{m}(\mathbf{x})\right)$
subject $\text{subject to}\quad\mathbf{x}\in\mathcal{X}$
Here, $\mathbf{x}\in\mathbb{R}^{n}$ denotes the decision variables, $\mathcal{X}$ the feasible domain, and $f_{i}(\mathbf{x})$ the $i$-th objective function ($i=1,2,\dots,m$). Due to trade-offs among the objectives, the goal is to find a set of Pareto optimal solutions rather than a single optimum.
Multi-objective optimization has been extensively studied in the context of path planning for over 50 years. Nevertheless, its application to object navigation (ON) remains largely unexplored. One reason for this may be that traditional ON settings typically assume a small-scale workspace (e.g., tens of meters square), where myopic and intuitive path planning is often sufficient. For example, these settings do not consider cases where landmark observations arrive incrementally or maps that grow without bounds, such as those commonly addressed in the SLAM community, including 1,000 km-scale autonomous driving scenarios [25] and vast open-sky 3D aerial drone scenarios [26].
In contrast, this study reformulates the standard single-target ON task as a multi-objective problem, called Multi-objective Optimization-driven Object Goal Navigation (MOON), with two conflicting objectives: (1) exploration of unvisited areas to discover unknown landmarks, and (2) exploitation of already mapped areas by efficiently revisiting previously-observed landmarks. The first objective corresponds to conventional ON methods, while the second objective leverages path optimization techniques, which is addressed for the first time in the current study. One implementation of the latter is detailed in Section III-A.
III-A MOON as a Variable-Horizon Set Orienteering Problem
We consider the following scenario: a robot plans a path over a navigation graph $G=(V,E)$ annotated with previously-observed landmarks. This graph only includes landmarks that have been observed within a radius $L$ from any of the robot’s previously visited viewpoints. That is, the navigation graph only contains landmarks the robot has visually observed, and only these landmarks are known to be visitable.
Under this constraint, the robot aims to plan a path from the currrent viewpoint that maximizes the total reward (i.e., landmark visits) within a time or distance budget. This can be formulated as a Set Orienteering Problem with Visibility Constraints.
  • Let $G=(V,E)$ be a directed graph, where $V$ is the set of nodes and $E$ the set of edges.
  • Each node $i\in V$ has an associated reward $p_{i}\geq 0$.
  • A start node $s\in V$ and an end node $t\in V$ are given.
  • Each edge $(i,j)\in E$ has an associated travel cost $c_{ij}\geq 0$.
  • A total travel budget $B>0$ is given.
  • Let $V^{\prime}\subseteq V$ be the set of landmarks observed within radius $L$ in the robot’s past trajectories (i.e., nodes included in the navigation graph). Only nodes in $V^{\prime}$ are eligible for visiting.
Variables
  • $x_{ij}\in\{0,1\}$: whether edge $(i,j)$ is included in the path.
  • $y_{i}\in\{0,1\}$: whether node $i$ is visited.
  • $u_{i}\in\mathbb{Z}$: auxiliary variable indicating visiting order (used to ensure path connectivity).
Objective
maximize: $\sum_{i\in V^{\prime}}p_{i}y_{i}$
Constraints
$\sum_{j\in V}x_{sj}$ $=1$ (1)
Start from node $s$
$\sum_{i\in V}x_{it}$ $=1$ (2)
End at node $t$ (optional if return is not required)
$\sum_{j\in V}x_{ji}$ $=y_{i}\quad\forall i\in V\setminus\{s\}$ (3)
In-degree consistency
$\sum_{j\in V}x_{ij}$ $=y_{i}\quad\forall i\in V\setminus\{t\}$ (4)
Out-degree consistency
$\sum_{(i,j)\in E}c_{ij}x_{ij}$ $\leq B$ (5)
Budget constraint
$u_{i}-u_{j}+1$ $\leq(|V|-1)(1-x_{ij})$ (6)
Subtour elimination: $\forall i\neq j,\;i,j\in V\setminus\{s,t\}$
$x_{ij}$$\in\{0,1\},\;y_{i}\in\{0,1\},\;u_{i}\in\mathbb{Z}\quad\forall i,j\in V$ (7)
III-B Extension to General-Purpose MOON
The proposed optimization approach for the Set Orienteering Problem (SOP) facilitates the seamless integration of established multi-objective optimization techniques. This integration is particularly pertinent in practical applications such as multi-objective navigation and robotic exploration & mapping, where objectives often extend beyond mere reward maximization to include minimizing travel distance, ensuring safety, enhancing information gain, and maximizing the probability of target detection. To effectively address these multifaceted requirements, it is advantageous to extend the SOP beyond a bi-objective framework (i.e., knowledge exploration and exploitation) and reformulate it as a general-purpose multi-objective optimization problem capable of simultaneously handling multiple evaluation criteria [27].
In multi-objectivizing the SOP, it is essential to first define multiple objective functions. In addition to the traditional reward function $r(v)$, other evaluation functions such as travel cost $c(v)$ and indicators for information gain or detection probability $p(v)$ can be introduced to enable richer assessments tailored to the task. Three main approaches are commonly used to integrate these objectives. First, scalarization assigns a weight $w_{i}$ to each objective and performs a linear combination to optimize a single scalar function. While this method is straightforward and computationally efficient, it may not fully capture the diversity of Pareto-optimal solutions [28]. Second, Pareto enumeration treats all objectives equally and seeks non-dominated solutions that are not inferior in any criterion. This approach offers greater flexibility but often incurs higher computational costs [29]. Third, the $\varepsilon$-constraint method focuses on optimizing one primary objective while treating others as constraints, providing practical controllability when specific priorities must be respected [30].
Effectively solving the multi-objective SOP in practice necessitates selecting an appropriate solver based on the problem’s structure and scale. Integer Linear Programming (ILP) approaches enable explicit formulation of multiple objectives through scalarization or $\varepsilon$-constraints, allowing for iterative optimization to enumerate multiple non-dominated solutions. Alternatively, metaheuristic methods such as Genetic Algorithms (GA) and NSGA-II are effective evolutionary approaches capable of exploring extensive solution spaces and approximating the Pareto front [29]. Furthermore, reinforcement learning-based approaches define the state in terms of current position, visit history, and remaining budget, selecting the next node to visit as an action. By designing value functions using a multi-head architecture, these approaches can simultaneously learn multiple objectives and make strategic decisions that adapt to situational demands [31, 32].
Figure 7: Navigation examples. This figure shows intermediate navigation results for three different workspaces. Observed layout elements, including rooms, walls, and doorways, are represented by black lines, while visited areas detected by the target detector are shown in gray. The workspace poses greater challenges to the agent than it might superficially appear, as the TFP agent lacks prior knowledge of room size, wall orientations, and the number and positions of doors. Training-free planners (TFP) are more suitable for our ablation study than learning-based planners such as RLP, since we aim to prevent ON performance from being influenced by prior knowledge.
IV Evaluation
IV-A Problem Setting
We consider a large-scale exploration problem in a workspace of size $300\,\mathrm{m}\times 300\,\mathrm{m}$, which is significantly larger than typical simulation environments such as Habitat-Sim. The workspace consists of multiple rooms, each having entrances and exits on four sides (see Fig. 7).
A single target point object is placed in the environment. While this work assumes a single target, it can be readily extended to a multi-target setting (multi-ON problem).
We assume a circular observation range centered at the robot’s viewpoint. This can be easily extended to a standard monocular camera with a field of view of around $40^{\circ}$.
IV-B Hyperparameters
IV-B1 User-defined Parameters
  • Long-range sensor radius $L=100,200\,\mathrm{m}$
  • Short-range sensor radius $R=3\,\mathrm{m}$
  • Number of landmarks $M=10,15,20$
  • Number of trials $U=$100
IV-B2 Randomized for Each Trial
  • Obstacle placement
  • Initial robot position
  • Target position
  • Landmark positions
IV-C Implementation Details
IV-C1 Object Grid Map
Before navigation begins, an object grid map is initialized as a blank map. During navigation, landmarks and obstacles (e.g., walls) detected within the long-range sensor radius $L$ are incrementally added to the map. Redundant entries are ignored, meaning that we assume zero information gain from repeated observations of the same landmark.
IV-C2 Graph Map Construction
  • Nodes: Each node is defined as a region of radius $R$ centered at an observed landmark. Previously observed landmarks are excluded from the navigation graph.
  • Edges: The cost between nodes is computed using the Dijkstra planner.
IV-C3 MOO Solver
The SOP solver from [33] is employed as the MOO solver of our system, and we have confirmed that it works efficiently with the navigation graph recognition module and the action planner module. This paper proposes a new Integer Linear Programming (ILP) formulation for small to medium-sized problems, alongside a Variable Neighborhood Search (VNS) metaheuristic-based algorithm (VNS-SOP) that helps reduce computation time. Beyond offering some improvements to existing benchmark results, it also demonstrates the applicability of its approach to other Orienteering Problem (OP) variants, such as the Dubins Orienteering Problem (DOP) and the Orienteering Problem with Neighborhoods (OPN), suggesting valuable insights.
IV-D Evaluation Metric
We use the Success weighted by Path Length (SPL) as the primary performance metric, defined as:
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\cdot\frac{l_{i}}{\max(p_{i},l_{i})}$ (8)
where $S_{i}\in\{0,1\}$ indicates whether the $i$-th trial was successful, $l_{i}$ is the shortest path length, and $p_{i}$ is the path length actually traveled.
IV-E Baselines
We compare our method against the following baselines:
  1. 1.
    Frontier-based exploration: A classical exploration strategy that does not utilize landmark information.
  2. 2.
    Sweeper robot planner
  3. 3.
    TSP-based planner
  4. 4.
    MOON-based planner
For the second planner, we experimented with multiple different implementations.
IV-F Results
In conclusion, both the second and third planners were found to be unsuitable for direct use in the current task, as they aim to sweep the entire floor area with the robot body and thus tend to generate overly dense paths. Of course, trying different spatial discretizations to generate graph nodes may still be effective and remains a topic for future work. The first and fourth planners, on the other hand, demonstrated meaningful SPL performance. The results are as follows:
  • L=100, R=3, M=10
    • Frontieer: 0.6083
    • MOON: 0.8315
  • L=100, R=3, M=15
    • Frontieer: 0.6395
    • MOON: 0.8581
  • L=100, R=3, M=20
    • Frontieer: 0.6473
    • MOON: 0.8734
  • L=100, R=10, M=10
    • Frontieer: 0.6540
    • MOON: 0.8395
  • L=100, R=10, M=15
    • Frontieer: 0.6373
    • MOON: 0.8188
  • LD100, R=10, M=20
    • Frontieer: 0.6425
    • MOON: 0.8309
The results indicate that the MOON planner consistently outperforms the Frontier planner.